OpenStudy (anonymous):

According to the commutative property of multiplication, which expression is equivalent to (3 + 7)(-2)? -2(3 + 7) 3(7 – 2) (3 + 7)(-2) 3 + 7 – 2 Which expression is equal to (21)(7)(3x) using the associative property? (21 · 7 · 3)x (7)(21)(3x) 32(7 + 3x) (3x)(7)(21) Use the distributive property to rewrite the expression 8(–2x – 3). –16x – 3 –16x + 3 –16x – 24 –16x + 24

OpenStudy (anonymous):

@BAdhi

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (badhi):

can I see your opinion on each of the questions please..

OpenStudy (anonymous):

B. for the first one

OpenStudy (anonymous):

C. for the thirrd one

OpenStudy (anonymous):

@BAdhi

OpenStudy (badhi):

commutative means if there is, (p+q+...)(c+d+...) -> we can write as (c+d+..)(p+q+...) for example (5+7)(3+4) is same as (3+4)(5+7) from the commutative property remember after application of any of these properties you should get the same result as the the expression that you began with so if we take your answer it give two different answers (3 + 7)(-2) = -20 where as 3(7 – 2) = 15

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

You have to use the distributive property.

OpenStudy (anonymous):

@BAdhi is my answer for the third one correct?

OpenStudy (badhi):

yes assosiative property means \((a\times b)\times (c) = (a)\times(b\times c) = (a\times b\times c)\)